# Flatness-based control of a Timoshenko beam

### 31 Mar 2026 v1

@@ I Introduction
This paper extends the flatness-based control design from [undefo] to hyperbolic MI systems with additional in-domain coupling. The proposed concept relies on the transformation into HCF, a system representation that is advantageous for feedback design. This is somewhat similar to the well-known backstepping method [undefh], which also employs a state transformation to simplify the control design. The main challenge of the HCF-based method lies in identifying a suitable candidate for the HCF state. As the flat output of systems with multiple inputs is not unique, the same applies to the HCF state [undefo]. Moreover, the HCF is no longer a state representation in the classical sense, in general, as it involves input predictions. Therefore, a non-static feedback of the HCF state is derived for trajectory tracking. It is shown that the transformation into HCF is invertible, allowing to express the control law in terms of the original system state.
@@ II Problem statement
The objective is to design a state feedback controller that stabilizes the transition of the Timoshenko beam (3)–(4) from an initial stationary configuration $(w(z,t_{0}),\,\varphi(z,t_{0}))=(w_{0}(z),\,\varphi_{0}(z))$ to a desired stationary configuration $(w(z,t_{T}),\,\varphi(z,t_{T}))=(w_{T}(z),\,\varphi_{T}(z))$ in finite time $T=t_{T}-t_{0}\in\mathbb{R}^{+}$. The control design proceeds in three steps:
@@ VI Simulation results
$\eta_{2}(\tau,t)$ $=\left\{\begin{aligned} &\bar{x}_{2}^{+}(\psi_{2}(\tau_{2}-\tau),t)-\textstyle\int_{\tau_{1}}^{\min(2\tau_{1},\tau_{1}+\tau)}b_{0,1}^{+}(\tau,\sigma)\bar{x}_{1}^{-}(\psi_{1}(\sigma-\tau_{1}),t)\,\mathrm{d}\sigma\\ &\hskip 18.49988pt-\textstyle\int_{\min(0,\tau-\tau_{1})}^{\tau-\tau_{1}}b_{0,2}^{+}(\tau,\sigma)\bar{u}_{1}(t+\sigma)\,\mathrm{d}\sigma,&&\qquad\tau\in[0,\tau_{2})\\ &-\bar{x}_{2}^{-}(\psi_{2}(\tau-\tau_{2}),t)-\textstyle\int_{\tau-\Delta\tau}^{\max(2\tau_{1},\tau-\Delta\tau)}b_{0,1}^{+}(\tau,\sigma)\bar{x}_{1}^{-}(\psi_{1}(\sigma-\tau_{1}),t)\,\mathrm{d}\sigma\\ &\hskip 18.49988pt-\textstyle\int_{\tau_{1}}^{\min(2\tau_{1},\tau-\Delta\tau)}b_{0,1}^{-}(\tau,\sigma)\bar{x}_{1}^{-}(\psi_{1}(\sigma-\tau_{1}),t)\,\mathrm{d}\sigma\\ &\hskip 18.49988pt-\textstyle\int_{\max(0,\tau-\tau_{2}-\tau_{1})}^{\tau-\tau_{1}}b_{0,2}^{+}(\tau,\sigma)\bar{u}_{1}(t+\sigma)\,\mathrm{d}\sigma-\textstyle\int_{\min(0,\tau-\tau_{2}-\tau_{1})}^{\tau-\tau_{2}-\tau_{1}}b_{0,2}^{-}(\tau,\sigma)\bar{u}_{1}(t+\sigma)\,\mathrm{d}\sigma,&&\qquad\tau\in[\tau_{2},2\tau_{2})\end{aligned}\right$. (30b)

### 01 Apr 2026 v2

@@ I Introduction
[…]
Figure 1: Schematic of the HCF-based control design.
This paper extends the flatness-based control design from [undefo] to hyperbolic MI systems with additional in-domain coupling. The proposed concept relies on the transformation into HCF, a system representation that is advantageous for feedback design (see Figure 1). This is somewhat similar to the well-known backstepping method [undefh], which also employs a state transformation to simplify the control design. The main challenge of the HCF-based method lies in identifying a suitable candidate for the HCF state. As the flat output of systems with multiple inputs is not unique, the same applies to the HCF state [undefo]. Moreover, the HCF is no longer a state representation in the classical sense, in general, as it involves input predictions. Therefore, a non-static feedback of the HCF state is derived for trajectory tracking. It is shown that the transformation into HCF is invertible, allowing to express the control law in terms of the original system state.
@@ II Problem statement
The objective is to design a state feedback controller that stabilizes the transition of the Timoshenko beam (3)–(4) from an initial stationary configuration $(w(z,t_{0}),\,\varphi(z,t_{0}))=(w_{0}(z),\,\varphi_{0}(z))$ to a desired stationary configuration $(w(z,t_{T}),\,\varphi(z,t_{T}))=(w_{T}(z),\,\varphi_{T}(z))$ in finite time $T=t_{T}-t_{0}\in\mathbb{R}^{+}$. The control design proceeds in three steps, as illustrated in Figure 1:
@@ VI Simulation results
$\eta_{2}(\tau,t)$ $=\left\{\begin{aligned} &\bar{x}_{2}^{+}(\psi_{2}(\tau_{2}-\tau),t)-\textstyle\int_{\tau_{1}}^{\min(2\tau_{1},\tau_{1}+\tau)}b_{0,1}^{+}(\tau,\sigma)\bar{x}_{1}^{-}(\psi_{1}(\sigma-\tau_{1}),t)\,\mathrm{d}\sigma\\ &\qquad-\textstyle\int_{\min(0,\tau-\tau_{1})}^{\tau-\tau_{1}}b_{0,2}^{+}(\tau,\sigma)\bar{u}_{1}(t+\sigma)\,\mathrm{d}\sigma,&&\quad\tau\in[0,\tau_{2})\\ &-\bar{x}_{2}^{-}(\psi_{2}(\tau-\tau_{2}),t)-\textstyle\int_{\tau-\Delta\tau}^{\max(2\tau_{1},\tau-\Delta\tau)}b_{0,1}^{+}(\tau,\sigma)\bar{x}_{1}^{-}(\psi_{1}(\sigma-\tau_{1}),t)\,\mathrm{d}\sigma\\ &\qquad-\textstyle\int_{\tau_{1}}^{\min(2\tau_{1},\tau-\Delta\tau)}b_{0,1}^{-}(\tau,\sigma)\bar{x}_{1}^{-}(\psi_{1}(\sigma-\tau_{1}),t)\,\mathrm{d}\sigma\\ &\qquad-\textstyle\int_{\max(0,\tau-\tau_{2}-\tau_{1})}^{\tau-\tau_{1}}b_{0,2}^{+}(\tau,\sigma)\bar{u}_{1}(t+\sigma)\,\mathrm{d}\sigma-\textstyle\int_{\min(0,\tau-\tau_{2}-\tau_{1})}^{\tau-\tau_{2}-\tau_{1}}b_{0,2}^{-}(\tau,\sigma)\bar{u}_{1}(t+\sigma)\,\mathrm{d}\sigma,&&\quad\tau\in[\tau_{2},2\tau_{2})\end{aligned}\right$. (30b)